# Revised Mass and Orbit of Eridani b: A 1 Jupiter-Mass Planet on a Near-Circular Orbit

### v1

@@ 2.4 Non-Linear Effects
In order to rigorously account for second order acceleration terms originating from the systemās close proximity to Earth, we model the three-dimensional motion of the systemās barycenter through space.
We convert the the proper motion, parallax, and radial velocity at a reference epoch, taken to be J2016, into Cartesian coordinates and propagate them linearly versus time assuming uniform space motion. This accounts for perspective (secular) acceleration, for curvature in the R.A. and Dec. spherical coordinate system, and for changing parallax versus time which causes an slight apparent magnification versus time effect.
In addition, we identify that that changing light travel-time is an important and previously neglected consideration for this systemāand perhaps many others. RV and pulsar timing data are often corrected for the so-called āRĆømer-delay,ā the changing light travel time between a source and the Earth due to the Earthās travel around the Sun. In addition to the Earthās motion, we note that the distance from our solar system to EridaniĀ has increased by 19 light-hours between 1980 and 2020. This has a small effect on the apparent orbital motion of the planet on the order of 1 mas, and on the perturbed stellar radial velocity on the order of 2 cm/s, but it has a significant effect on the observed long-term proper motion: 975 mas/yr 19 hr = 2.1 mas less displacement than expected from straight-line motion extrapolated over this time span. This is larger than the astrometric uncertainties of missions like Gaia and Hipparcos.
In the context of the Hipparcos-Gaia position difference, the apparent time delay of 11.5 hr corresponds to an apparent long term proper motion that is 0.05 mas/yr lower than a naive model would suggest, an effect greater than the uncertainty reported by the HGCA. Precision orbit models of high proper motion stars should take this effect into account when modeling long-term proper motion.
To account for this changing light travel time in our models, we solve for the emission time at each epoch that corresponds to our observation time using an iterative scheme. We calculate this difference compared to the reference epoch J2016, for which we set ; that is, we donāt consider the total light travel time, only the change in light travel time from the systemās distance at the reference epoch.
Lastly, we do not consider the third order perturbations to this motion caused by the planet; that is, the above effects are calculated for the systemās barycenter applied to both the star and planet.
We make these corrections at every epoch and model evaluation since the barycentric proper motion of the system is allowed to vary, and projecting the (near-) instantaneous proper motion from Gaia back would unintionally mix in contributions from the planetās orbit.
@@ 2.5 Gaussian Process
EridaniĀ has significant photometric and RV variability consistent with a rotation period of 10-12 days. As with the works of Llop-Sayson etĀ al. (2021), Giguere etĀ al. (2016), and Roettenbacher etĀ al. (2022), we model this activity signal using a Gaussian process with a quasi-periodic kernel. We implemented the GP in our model using the Celerite.jl package from Foreman-Mackey etĀ al. (2017), after forking and updating the package to work in modern Julia. We fit the GP separately to each dataset, but share a single set of GP hyper-parameters between all instruments.
6em
@@ 2.6 Implementation
We implemented our models in the Julia-based (Bezanson etĀ al., 2012) Octofitter framework (Thompson etĀ al., 2023) and sample using stabilized variational non-reversible parallel tempering Surjanovic etĀ al. (2022); Syed etĀ al. (2021) implemented in Pigeons.jl (Surjanovic etĀ al., 2023).
Unlike the model presented in Llop-Sayson etĀ al. (2021), and like the models presented in Feng etĀ al. (2023) and FritzĀ Benedict (2022) we jointly model both the astrometry and radial velocity, rather than model the RV and astrometry separately. Unlike Feng etĀ al. (2023) and FritzĀ Benedict (2022), we include the Gaussian process model of stellar activity. As with FritzĀ Benedict (2022) we include the Hubble FGS astrometry originally published in Benedict etĀ al. (2006).
The largest model which incorporates every listed data source has a total of 61 different parameters. The majority of these are instrument-specific nuisance parameters, which we marginalize over when presenting results. The full list of parameters and our adopted priors are listed in Table 1. Whenever we sample from a model that ignores a certain dataset, we remove the related instrument-specific nuisance parameters (e.g. a model without any RV data does not include the many RV offsets, jitters, and slope parameters).
We sample each model until the log-evidence converges between sampling rounds and the diagnostic (Gelman & Rubin, 1992; Vehtari etĀ al., 2021) is to at least two or more decimal places.
@@
APF zero point | m/s | ||
EXPRES zero point | m/s | ||
CHIRON zero point | m/s | ||
CFHT zero point | m/s | ||
HIRES zero point | m/s | ||
Lick (1) zero point | m/s | ||
Lick (2) zero point | m/s | ||
Lick (3) zero point | m/s | ||
Lick (4) zero point | m/s | ||
HARPS (pre) zero point | m/s | ||
HARPS (post) zero point | m/s | ||
CES (lc) zero point | m/s | ||
CES (vlc) zero point | m/s | ||
@@ 2.7 List of Models
In order to assess the agreement between different datasets and rule out systematic errors, we create a set of 7 different models. We progressively add additional data to build up to the complete model. We compare the parameters inferred by each model and their predictions for the planetās location to study what constraints come from which dataset. Checking that the inferred parameters and predictions are not in tension between models will allow us to check for signs of systematic errors that could bias our results.
@@
Full model with RV, HGCA, FGS, and the DR vs. DR3 acceleration.
The 1st (RV only) and 2nd (astrometry only) models include entirely separate data, aside from the position and parallax of the star from Gaia DR3. Similarly, if we consider that the orbital semi-major axis, eccentricity, and argument of periastron are determined largely by RV data, we can then also compare the nearly independent constraints on the inclination and position angle of ascending node from the Hipparcos IAD, the Gaia-Hipparcos acceleration, and the Hubble FGS data using results from the 3rd, 4th, and 5th models.
-1.45cm-1.2cm
-1.8cm-0.3cm
@@ 3 Results
We now present the results of our models, starting with only the RV data and building up to the final solution. The posteriors from all seven models are compared in Figure 1 and 2, which includes a corner plot and predictions for 2025-01-01. We use consistent colors to identify each model throughout the figures in this section. We report results using a mix of credible intervals (CI) and highest density intervals (HDI) as appropriate to summarize peaked or truncated distributions, with the notation used to indicate the median and 14% / 84% quantiles, and the [lower,upper] notation to indicate the the 75% HDI interval. As with other works, we use 75% to emphasize the arbitrariness of the specific probability mass fraction. Tables summarizing the posteriors of each model are presented in the Appendix.
@@ 3.2 Constraints added by the Hipparcos IAD
Adding the Hipparcos IAD data provides constraints on the inclination, position angle of ascending node, and position angle of the planet versus time. This constrains the planet to be roughly on the South side of the star in 2025, North in 2029, and so on. It provides essentially no information on the horizontal position of the planet. The addition of the Hipparcos IAD does not significantly constrain the mass of the planet, other than that it is less than 1.3 Ā (84 percentile upper limit), or the 75% HDI of [0.56, 1.09] .
In the versus subpanel of Figure 2, we see that the Hipparcos IAD data does not provide any strong constraints on the inclination of the planetās orbit, but it does suggest that if the planet is co-planar with the outer disk, then it is more likely to orbiting counterclockwise.
@@ 3.3 Constraints added by the HGCA
Adding the DR3 Hipparcos-Gaia proper motion anomaly from the eDR3 HGCA (Brandt, 2021) to the RV data provides significantly stronger constraints. Note that this model includes only net velocities from the Gaia and Hipparcos cataloges, and does not include the within-catalogue motion from the Hipparcos IAD. With RV and the HGCA data, we find a mass upper limit (84th percentile) of 1.00 . With regard to the orientation of the orbit, the HGCA data constrains it to a bimodal distribution of either or , and a position angle of ascending node of , both of which are in agreement with the orientation of the outer debris disk to (Booth etĀ al., 2023).
@@
Figure 4 shows the RV, proper motion, and position perturbation in R.A. and Dec. versus time for different orbit draws. Looking at the proper motion panels, the orbit is near its most extreme values in the Dec. direction, but at mid-phase in R.A.āhence why the motion in Dec. is well-determined but ambiguous in R.A. with these data alone.
@@ 3.4 Constraints added by the Hubble FGS data
Adding the Hubble FGS data from Benedict etĀ al. (2006) independently constrains the planet to be SSE or S of the star in 2025. These data prefer a slightly lower 84% upper mass limit of 0.89 .
@@ 3.5 Constraints added by the Gaia DR3-DR2 acceleration
Similarly to the addition of the FGS data, adding the DR3-DR2 acceleration removes an ambiguity in the R.A. motion of the star. Notably, the constraint provided by either the DR3-DR2 acceleration or the Hubble FGS astrometry selects the same part of the ambiguous posterior from the RV & HGCA alone, increasing our confidence in the model.
@@ 3.6 Complete Model
With the complete model, the planet is very well-localized (Figure 2, top right panel). The model yields a semi-major axis of AU, a mass of , eccentricity of [0.00,0.09], and an orbit that is inclined . The direction of motion is counterclockwise in the plane of the sky.
@@ 3.7 Astrometry only orbit solution
The data from Hipparcos, Hubble FGS, Gaia DR2, and DR3 is sufficient to, for the first time, fit an astrometry-only model, without any RV data. Using a uniform prior on the orbital period of 1ā10 yr, we find an orbit posterior with a semi-major axis of AU. We find the model prefers moderately low eccentricity solutions (HDI of [0, 0.42], peaked at 0). The orbital period agrees to better than with the RV only model. The times of conjunction and opposition are likewise moderately well constrained, though not as precisely as by the RV data. The astrometry data independently constrains the mass of the planet to .
@@ 4.1 Orbital Period
All of our orbit models that include RV data support an orbital period in the range of 7.3ā7.45 years, with the complete model giving yr. This is very similar to the period of yr reported by Llop-Sayson etĀ al. (2021), though somewhat shorter than the yr or yr periods reported by Feng etĀ al. (2023) and by Benedict etĀ al. (2006) respectively. We believe this disagreement stems from differences in the noise model used to deal with stellar activity, which we discuss in Section 4.3.
Our model of the absolute astrometry only, with no RV data, supports a period of years. This independent support for the orbital period should lay to rest concerns that the RV signature of Eridani bĀ may be an artifact of long-term stellar activity cycles.
Indeed, the perturbation orbit of the star has a projected semi-major axis of 2.7 mas which is just over twice the 1.1 mas apparent radius of the star (using a 0.738 stellar radius from Rains etĀ al. (2020)). Such a large displacement of the photocenter over a period of years cannot be explained by stellar activity.
@@ 4.2 Orbital Phase
The RV model we present determines the overall phase of the orbit and constrains the times of conjunction and opposition to approximately one month. Given the known parallax of the system, this means that the RV data alone can constrain the separation of the planet to ring whenever it is near oppositionāin 2017, 2021, 2025, 2029, and so on. This was noted as well in Mawet etĀ al. (2019).
The work by Llop-Sayson etĀ al. (2021) added constraints from Hipparcos and Gaia DR2 and presented predictions for the planetās location at two imaging epochs. The astrometry data added constraints to the PA versus time, but the additional data broadened the constraints to the planetās separation compared to RV data alone. We hypothesize this may be a result of the two-step process in which the RV data were modeled first, a kernel density estimate (KDE) was fit to the posterior, and then the KDE was used as a prior in the subsequent absolute astrometry fit. The variables used in the astrometry fit do not include the time of conjunction directly, expressing it instead as a combination of other variables. It could be that while the KDE accurately represented the marginal posteriors of the period, eccentricity, argument of periastron, and epoch of periastron passage, the detailed correlation structure may not have been captured with sufficient detail to give a tightly constrained prior on time of conjunction derived from a combination of the above.
Joint models like the ones presented here and in previous works Benedict etĀ al. (2006); Reffert & Quirrenbach (2011); FritzĀ Benedict (2022); Feng etĀ al. (2023) sidestep this challenge. This could explain why our RV and Hipparcos model has considerably less uncertainty on the separation of the planet near the times of conjunction (e.g. in 2017-09-01, see Figure 9).
6em
@@ 4.3 Eccentricity
The eccentricity of the planet is a key concern since it has a strong impact its separation from the star at opposition. Our models all support a low eccentricity that is consistent with zero. This is consistent with previous works that included a GP model of stellar activity (Mawet etĀ al., 2019; Llop-Sayson etĀ al., 2021; Giguere etĀ al., 2016) (eccentricity near zero) and is inconsistent with works that did not such as Feng etĀ al. (2023) and FritzĀ Benedict (2022) (eccentricity of 0.26 and 0.16 respectively).
6em

### v2

@@ 2.4 Non-Linear Effects
In order to rigorously account for second order acceleration terms originating from the system’s close proximity to Earth, we model the three-dimensional motion of the system’s barycenter through space.
We convert the proper motion, parallax, and radial velocity at a reference epoch, taken to be J2016, into Cartesian coordinates and propagate them linearly versus time assuming uniform space motion. This accounts for perspective (secular) acceleration, for curvature in the R.A. and Dec. spherical coordinate system, and for changing parallax versus time which causes an slight apparent magnification versus time effect.
In addition, we include the changing light travel-time to the system in our model. RV and pulsar timing data are often corrected for the so-called “Rømer-delay,” the changing light travel time between a source and the Earth due to the Earth’s travel around the Sun. In addition to the Earth’s motion, we note that the distance from our solar system to Eridani has increased by 19 light-hours between 1980 and 2020. This has a small effect on the apparent orbital motion of the planet on the order of 1 mas, and on the perturbed stellar radial velocity on the order of 2 cm/s. It also has an impact of the observed long-term proper motion on the order of 2 mas less displacement than expected from straight-line motion extrapolated over this time span.
To account for this changing light travel time in our models, we solve for the emission time at each epoch that corresponds to our observation time using an iterative scheme. We calculate this difference compared to the reference epoch J2016, for which we set ; that is, we don’t consider the total light travel time, only the change in light travel time from the system’s distance at the reference epoch.
We make these corrections at every epoch and model evaluation since the barycentric proper motion of the system is allowed to vary, and projecting the (near-) instantaneous proper motion from Gaia back would unintentionally mix in small contributions from the planet’s orbit.
We do not consider the third order perturbations to this motion caused by the planet; that is, the above effects are calculated for the system’s barycenter applied to both the star and planet.
Although we include the light travel time effect in our model, for typical stars and timescales it is very well approximated as a linear function appearing to increase the proper motion measured by both Hipparcos and Gaia. As such, readers should note that our proper motion parameters do not include this term and are therefore approximately 0.05 mas/yr lower than the apparent sky motion. Future orbit models of precision data, e.g. combining Gaia DR4 and high precision interferometry should take this effect into account when modeling long-term proper motion.
@@ 2.5 Gaussian Process
Eridani has significant photometric and RV variability consistent with a rotation period of 10-12 days. As with the works of Llop-Sayson et al. (2021), Giguere et al. (2016), and Roettenbacher et al. (2022), we model this activity signal using a Gaussian process with a quasi-periodic kernel. We implemented the GP in our model using the Celerite.jl package from Foreman-Mackey et al. (2017), after forking and updating the package to work in modern Julia. We fit the GP separately to each dataset, but share a single set of GP hyper-parameters between all instruments.
@@ 2.6 Implementation
We implemented our models in the Julia-based (Bezanson et al., 2012) Octofitter framework (Thompson et al., 2023) and sample using stabilized variational non-reversible parallel tempering Surjanovic et al. (2022); Syed et al. (2021) implemented in Pigeons.jl (Surjanovic et al., 2023).
Unlike the model presented in Llop-Sayson et al. (2021), and like the models presented in Feng et al. (2023) and Fritz Benedict (2022) we jointly model both the astrometry and radial velocity, rather than model the RV and astrometry separately. Unlike Feng et al. (2023) and Fritz Benedict (2022), we include the Gaussian process model of stellar activity. As with Fritz Benedict (2022) we include the Hubble FGS astrometry originally published in Benedict et al. (2006).
The most complete model which incorporates every listed data source has a total of 66 different parameters. The majority of these are instrument-specific nuisance parameters, which we marginalize over when presenting results. The full list of parameters and our adopted priors are listed in Table 1. Whenever we sample from a model that ignores a certain dataset, we remove the related instrument-specific nuisance parameters (e.g. a model without any RV data does not include the many RV offsets, jitters, and slope parameters).
We sample each model until the log-evidence converges between sampling rounds and the diagnostic (Gelman & Rubin, 1992; Vehtari et al., 2021) is to at least two or more decimal places.
@@
| APF zero point | m/s | Uniform(-1000, 1000) | |
| EXPRES zero point | m/s | Uniform(-1000, 1000) | |
| CHIRON zero point | m/s | Uniform(-1000, 1000) | |
| CFHT zero point | m/s | Uniform(-1000, 1000) | |
| HIRES zero point | m/s | Uniform(-1000, 1000) | |
| Lick (1) zero point | m/s | Uniform(-1000, 1000) | |
| Lick (2) zero point | m/s | Uniform(-1000, 1000) | |
| Lick (3) zero point | m/s | Uniform(-1000, 1000) | |
| Lick (4) zero point | m/s | Uniform(-1000, 1000) | |
| HARPS (pre) zero point | m/s | Uniform(-1000, 1000) | |
| HARPS (post) zero point | m/s | Uniform(-1000, 1000) | |
| CES (lc) zero point | m/s | Uniform(-1000, 1000) | |
| CES (vlc) zero point | m/s | Uniform(-1000, 1000) | |
@@ 2.7 List of Models
In order to assess the agreement between different datasets and rule out systematic errors, we create a set of 7 different models. We progressively add additional data to build up to the complete model. We compare the parameters inferred by each model and their predictions for the planet’s location to study what constraints come from which dataset. Checking that the inferred parameters and predictions are not in tension between models will allow us to check for signs of systematic errors that could bias our results.
@@
Full model with RV, HGCA, FGS, and the DR2 vs. DR3 acceleration.
The 1st (RV only) and 2nd (astrometry only) models include entirely separate data, aside from the position and parallax of the star from Gaia DR3. Similarly, if we consider that the orbital semi-major axis, eccentricity, and argument of periastron are determined largely by RV data, we can then also compare the nearly independent constraints on the inclination and position angle of ascending node from the Hipparcos IAD, the Gaia-Hipparcos acceleration, and the Hubble FGS data using results from the 3rd, 4th, and 5th models. Note that the Hipparcos IAD residuals are only used in the 3rd model—the others use the calibrated Hipparcos catalog parameters from the HGCA and only use the measurement epochs from the IAD.
@@ 3 Results
We now present the results of our models, starting with only the RV data and building up to the final solution. The posteriors from all seven models are compared in Figure 1 and 2, which includes a corner plot and predictions for 2025-01-01. We use consistent colors to identify each model throughout the figures in this section. We report results using a mix of credible intervals (CI) and highest density intervals (HDI) as appropriate to summarize peaked or truncated distributions, with the notation used to indicate the median and 14% / 84% quantiles, and the [lower,upper] notation to indicate the the 75% HDI interval. As with other works, we use 75% to emphasize the arbitrariness of the specific probability mass fraction. Tables summarizing the posteriors of each model are presented in the Appendix. The final posterior corresponding to model 7 is available by request.
@@ 3.2 Constraints added by the Hipparcos IAD
Adding the Hipparcos IAD data provides constraints on the inclination, position angle of ascending node, and position angle of the planet versus time. This constrains the planet to be roughly on the South side of the star in 2025, North in 2029, and so on. It provides essentially no information on the horizontal position of the planet. The addition of the Hipparcos IAD does not significantly constrain the mass of the planet, other than that it is less than 1.3 (84 percentile upper limit), or the 75% HDI of [0.56, 1.09] .
In the versus subpanel of Figure 2, we see that the Hipparcos IAD data does not provide any strong constraints on the inclination of the planet’s orbit, but it does suggest that if the planet is co-planar with the outer disk, then it is more likely to orbiting counterclockwise.
@@ 3.3 Constraints added by the HGCA
Adding the DR3 Hipparcos-Gaia proper motion anomaly from the eDR3 HGCA (Brandt, 2021) to the RV data provides significantly stronger constraints. Note that this model includes only net velocities from the Gaia and Hipparcos cataloges, and does not include the within-catalogue motion from the Hipparcos IAD. With RV and the HGCA data, we find a mass upper limit (84th percentile) of 1.00 . With regard to the orientation of the orbit, the HGCA data constrains the inclination to a bimodal distribution of either or , and a position angle of ascending node of , both of which are in agreement with the orientation of the outer debris disk to (Booth et al., 2023).
@@
Figure 4 shows the RV, proper motion, and position perturbation in R.A. and Dec. versus time for different orbit draws. Looking at the proper motion panels, the orbit is near its most extreme values in the Dec. direction, but at mid-phase in R.A.—hence why the motion in Dec. is well-determined but ambiguous in R.A. with these data alone.
@@ 3.4 Constraints added by the Hubble FGS data
Adding the Hubble FGS data from Benedict et al. (2006) independently constrains the planet to be SSE or S of the star in 2025. These data prefer a slightly lower 84% upper mass limit of 0.89 .
@@ 3.5 Constraints added by the Gaia DR3-DR2 acceleration
The addition of the DR2 data acts to significantly remove the ambiguity in the otherwise bimodal inclination marginal inclination posterior from the RV, Hipparcos, and DR3 model (employing the HGCA). The result is similar to the addition of the FGS data, in that it acts to removes an ambiguity in the R.A. motion of the star. Notably, the constraint provided by either the DR3-DR2 acceleration or the Hubble FGS astrometry selects the same part of the ambiguous posterior from the RV & HGCA alone, increasing our confidence in the model.
@@ 3.6 Complete Model
With the complete model, the planet is very well-localized (Figure 2, top right panel). The model yields a semi-major axis of AU, a mass of , eccentricity of [0.00,0.10], and an orbit that is inclined . The direction of motion is counterclockwise in the plane of the sky.
@@ 3.7 Astrometry only orbit solution
The data from Hipparcos, Hubble FGS, Gaia DR2, and DR3 is sufficient to, for the first time, fit an astrometry-only model, without any RV data. Using a uniform prior on the orbital period of 0.5–100 yr, we find an orbit posterior that is multi-modal with two prominent modes, one at 7.2 yr and one at 17 yr. The 7.2 yr mode gives a constrained orbit that agrees very well with the RV only solution: consistent eccentricity, time of conjunction, , and low to moderate eccentricity (HDI of [0, 0.42], peaked at 0). This mode is also consistent with the orientation of the outer disk. The 17 yr mode on the other hand is inconsistent with the RV data. It has a higher and is inconsistent with the orientation of the outer disk. We further experimented with a full joint RV and astrometry model with period forced to be near the 17 yr mode identified from the astrometry. This resulted in a high eccentricity solution that was significantly disfavored versus the unconstrained RV and astrometry model according to the Bayes factor. We therefore perform a broad cut of the posterior to select select the 7 yr model by adopting a uniform prior on the period of 5.36 to 9.36 yr (see Table 1). The orbital period from the remaining astrometry only samples agrees within with the RV only model. The times of conjunction and opposition are likewise moderately well constrained and in agreement. The astrometry data provides a near-independent constraint on the mass of the planet to , with the caveat of us selecting the mode that that agrees with the orientation of the disk and the RV period.
@@ 4.1 Orbital Period
All of our orbit models that include RV data support an orbital period in the range of 7.3–7.45 years, with the complete model giving yr. This is very similar to the period of yr reported by Llop-Sayson et al. (2021), though somewhat shorter than the yr or yr periods reported by Feng et al. (2023) and by Benedict et al. (2006) respectively. We believe this disagreement stems from differences in the noise model used to deal with stellar activity, which we discuss in Section 4.3.
Our model of the absolute astrometry only, with no RV data, supports a period of years. This independent support for the orbital period should lay to rest concerns that the RV signature of Eridani b may be an artifact of long-term stellar activity cycles.
Indeed, the perturbation orbit of the star has a projected semi-major axis of 2.7 mas which is just over twice the 1.1 mas apparent radius of the star (using a 0.738 stellar radius from Rains et al. (2020)). Such a large displacement of the photocenter over a period of years cannot be explained by stellar activity.
@@ 4.2 Orbital Phase
The RV model we present determines the overall phase of the orbit and constrains the times of conjunction and opposition to approximately one month. Given the known parallax of the system, this means that the RV data alone can constrain the separation of the planet to ring whenever it is near opposition—in 2017, 2021, 2025, 2029, and so on. This was noted as well in Mawet et al. (2019).
The work by Llop-Sayson et al. (2021) added constraints from Hipparcos and Gaia DR2 and presented predictions for the planet’s location at two imaging epochs. The astrometry data added constraints to the PA versus time, but the additional data broadened the constraints to the planet’s separation compared to RV data alone. We hypothesize this may be a result of the two-step process in which the RV data were modeled first, a kernel density estimate (KDE) was fit to the posterior, and then the KDE was used as a prior in the subsequent absolute astrometry fit. The variables used in the astrometry fit do not include the time of conjunction directly, expressing it instead as a combination of other variables. It could be that while the KDE accurately represented the marginal posteriors of the period, eccentricity, argument of periastron, and epoch of periastron passage, the detailed correlation structure may not have been captured with sufficient detail to give a tightly constrained prior on time of conjunction derived from a combination of the above.
Joint models like the ones presented here and in previous works Benedict et al. (2006); Reffert & Quirrenbach (2011); Fritz Benedict (2022); Feng et al. (2023) sidestep this challenge. This could explain why our RV and Hipparcos model has considerably less uncertainty on the separation of the planet near the times of conjunction (e.g. in 2017-09-01, see Figure 9).
@@ 4.3 Eccentricity
The eccentricity of the planet is a key concern since it has a strong impact its separation from the star at opposition. Our models all support a low eccentricity that is consistent with zero. This is consistent with previous works that included a GP model of stellar activity (Mawet et al., 2019; Llop-Sayson et al., 2021; Giguere et al., 2016) (eccentricity near zero) and is inconsistent with works that did not such as Feng et al. (2023) and Fritz Benedict (2022) (eccentricity of 0.26 and 0.16 respectively).